Locate the absolute maximum and minimum for each of the following functions and justify your responses. (Show all work)

Question 1: 𝑓(π‘₯) = Cube root of x^2 βˆ’ π‘₯ on the interval [βˆ’1,1]Β 

Question 2: g(x)=xe^2x on the interval [-2,0]



First and Second Derivative AnalysisΒ :For each of the following functions, respond to the given prompts. Show all work that leads to your responses:


Given 𝑓(π‘₯) = 3π‘₯ 3 βˆ’ 18π‘₯ 2 βˆ’ 45π‘₯ + 10


a. On what interval(s), is 𝑓(π‘₯) increasing? Justify. show all work


b. At what value(s) of π‘₯ does 𝑓(π‘₯) have a relative minimum? Justify. show all work


c. On what interval(s), is 𝑓(π‘₯) decreasing and concave up? Justify. show all work




1
Expert's answer
2021-02-02T05:08:27-0500

The absolute maximum and minimum

"\mathrm{Suppose\:that\:}x=c\mathrm{\:is\:a\:critical\:point\:of\:}f\left(x\right)\mathrm{\:then,\:}"

"\mathrm{If\:}f\:'\left(x\right)>0\mathrm{\:to\:the\:left\:of\:}x=c\mathrm{\:and\:}f\:'\left(x\right)<0\mathrm{\:to\:the\:right\:of\:}x=c\mathrm{\:then\:}x=c\mathrm{\:is\:a\:local\:maximum.}"

"\mathrm{If\:}f\:'\left(x\right)<0\mathrm{\:to\:the\:left\:of\:}x=c\mathrm{\:and\:}f\:'\left(x\right)>\:0\mathrm{\:to\:the\:right\:of\:}x=c\mathrm{\:then\:}x=c\mathrm{\:is\:a\:local\:minimum.}"

"\mathrm{If\:}f\:'\left(x\right)\mathrm{\:is\:the\:same\:sign\:on\:both\:sides\:of\:}x=c\mathrm{\:then\:}x=c\mathrm{\:is\:neither\:a\:local\:maximum\:nor\:a\:local\:minimum.}"

"f'(x)" gives stationary points. Hence differentiation with respect to x


1) Applying chain rule "f'(x)=\frac{1}{3\left(x^2-x\right)^{\frac{2}{3}}}\frac{d}{dx}\left(x^2-x\right)=\frac{1}{3\left(x^2-x\right)^{\frac{2}{3}}}\left(2x-1\right)"

"f'(x)=\frac{2x-1}{3\left(x^2-x\right)^{\frac{2}{3}}}"

Hence "(x,f(x))=(−1,-2\sqrt3)" for absolute maximum

Hence "(x,f(x))=(−1,-\frac{\sqrt3}{4})" for absolute minimum


2) First derivative for "3x^3−18x^2−\:45x+10"

"\implies f'(x)=\frac{d}{dx}\left(3x^3\right)-\frac{d}{dx}\left(18x^2\right)-\frac{d}{dx}\left(45x\right)+\frac{d}{dx}\left(10\right)"

"\implies f'(x)=9x^2-36x-45"

Second derivative for "3x^3−18x^2−\:45x+10"

"\implies f''(x)=\frac{d}{dx}(9x^2)-\frac{d}{dx}(36x)-\frac{d}{dx}(45)"

"\implies f''(x)=18x-36"


a. Increasing interval(s) of 𝑓(π‘₯) "\mathrm{If\:}f\:'\left(x\right)>0\mathrm{\:to\:the\:left\:of\:}x=c\mathrm{\:and\:}f\:'\left(x\right)<0\mathrm{\:to\:the\:right\:of\:}x=c\mathrm{\:then\:}x=c\mathrm{\:is\:a\:local\:maximum.}"

Hence "f'(x)>0" is the increasing intervals "\implies" "(-\infty <x<-1) and (5<x<\infty )"


b. A relative minimum is when "f'\left(x\right)=9x^2−36x−45=0"

"x_{1,\:2}=\frac{-\left(-36\right)\pm \sqrt{\left(-36\right)^2-4\cdot \:9\left(-45\right)}}{2\cdot \:9}"

"x_{1,\:2}=\frac{-\left(-36\right)\pm \:54}{2\cdot \:9}"

"x_1=\frac{-\left(-36\right)+54}{2\cdot \:9},\:x_2=\frac{-\left(-36\right)-54}{2\cdot \:9}"

"x=5,\:x=-1"

"\mathrm{Minimum}\left(5,\:-290\right)"

At x=5, 𝑓(π‘₯) have a relative minimum


c. Decreasing interval(s) of 𝑓(π‘₯) "\mathrm{If\:}f\:'\left(x\right)<0\mathrm{\:to\:the\:left\:of\:}x=c\mathrm{\:and\:}f\:'\left(x\right)>0\mathrm{\:to\:the\:right\:of\:}x=c\mathrm{\:then\:}x=c\mathrm{\:is\:a\:local\:maximum.}"

Hence "f'(x)<0" is the decreasing interval "\implies (-1<x<5)"